# MoVieS: Motion-Aware 4D Dynamic View Synthesis in One Second

### v1

@@ 4.4 Ablation and Analysis
We investigate different camera conditioning strategies in the static pertaining stage. Quantitative comparisons are provided in Table 5. Camera tokens are injected throughout the feature backbone, enabling effective camera-aware modeling. In contrast, Plücker embeddings provide limited conditioning on their own and are merely comparable to having no camera information at all. However, as a pixel-aligned representation, Plücker embeddings are complementary to camera tokens, and their combination yields the most effective camera conditioning.
@@
To learn object 3D movements in dynamic scenes, we provide two kinds of motion supervision: (1) point-wise L1 loss and (2) distribution loss, as described in Equation 4. Their efficiency is evaluated via the 3D point tracking task in Table 5. Without any motion supervision, i.e., learning solely from novel view synthesis, training exhibits severe loss oscillations and frequent None gradients. The distribution loss captures only relative motion between pixels, while point-wise L1 loss produces more reasonable motion maps. Combining both leads to sharper boundaries. Qualitative results of estimated motions from different motion objectives are provided in Figure 4.
@@
Thanks to the unified design of MoVieS, it supports simultaneous novel view synthesis (NVS) and motion estimation. We study their mutual benefits in Table 6. “NVS w/o motion” disables explicit motion supervision during training, relying solely on NVS as a proxy to learn dynamics. As shown in Table 6 and Figure 4, this setting fails to learn meaningful motion and tends to model static scenes. “Motion w/o NVS” detaches the motion head from 3DGS rendering and instead conditions the depth head on time. Although explicit supervision enables some motion learning, the predictions are blurry and low-quality, as shown in Figure 4. Moreover, the depth head must now model both geometry and dynamics, increasing its burden and negatively affecting NVS. These results highlight the mutual reinforcement between NVS and motion estimation in MoVieS, where joint training leads to better performance on both.
@@ 4.5 Zero-shot Applications
Scene flow can be naturally derived by transforming the estimated per-pixel motion vectors from world coordinates to the target camera’s coordinates. Visualization results in Figure 5(a) present sharp edges and accurate motion directions. More results are provided in Figure 8.
@@
By thresholding the norm of per-pixel motion vectors, the estimated motion maps can be used to segment moving objects (Figure 5(b)), which is an essential task in computer vision and robotics (xie2024moving, ; huang2025segment, ). Remarkably, this is achieved without any explicit mask supervision, demonstrating the strong potential of our method. More results are provided in Figure 9.
Figure 5: Zero-shot Applications. The predicted pixel-aligned motion maps from our model can be directly applied to downstream tasks, such as (a) scene flow estimation and (b) moving object segmentation, in a zero-shot manner, without any task-specific fine-tuning or supervision.
@@ 5 Conclusion
In this work, we introduce MoVieS, a feed-forward model for dynamic novel view synthesis from monocular videos. It’s trained on large-scale datasets from diverse sources and jointly models scene appearance, geometry and motion in a unified and efficient network. Dynamic splatter pixels are proposed to represent dynamic scenes, enabling accurate and temporally coherent 4D reconstruction. Beyond novel view synthesis, MoVieS supports a broad range of applications including depth estimation, 3D point tracking, scene flow estimation and moving object segmentation, demonstrating its versatility for dynamic scene perception. We hope this work could serve as a step toward generalizable dynamic scene understanding, and support applications that require spatial and motion intelligence. More discussions on limitations, future work and broader impact are included in Appendix B.
@@ A.1 Dataset Details
As summarized in Table 1, the majority of training datasets used in our experiments are synthetic, providing rich annotations such as pixel-aligned depth maps and accurate camera intrinsics and extrinsics. Ground-truth 3D point trajectories are also available in PointOdyssey [33] and DynamicReplica [34]. Stereo4D [35], on the other hand, is a large-scale dataset constructed from YouTube stereo videos and annotated using pretrained foundation models [62, 98]. Despite being partially noisy, its diversity and scale offer strong generalization benefits.
The utilized backbone VGGT [10] requires all 3D scenes to be normalized to a unit scale, which in turn necessitates depth maps and camera extrinsics to be defined in a consistent metric space. This requirement is satisfied by the synthetic datasets and Stereo4D, whose camera poses and depths are metric-aligned by construction. In contrast, RealEstate10K [28] only provides relative camera parameters estimated via COLMAP [6], resulting in an unknown global scale. To address this, we re-estimate both depth maps and camera extrinsics using recent foundation models, Video Depth Anything [5] and MegaSaM [101], to recover aligned geometry across frames.
@@ A.2 Curriculum Training
We observed that training MoVieS is particularly unstable: the training loss often fluctuates abruptly, and gradients are prone to becoming None. It may arise from sparse annotations and the heterogeneous nature of the training data, which mixes datasets from various sources with differing domains (e.g., indoor vs. outdoor, real vs. synthetic), camera fields of view, recording frame rates, etc.
Thanks to the versatile design of MoVieS, we employ a curriculum strategy that gradually increases the training complexity. It begins by pretraining the model on low-resolution ($224\times 224$) static datasets with only depth and photometric losses, then introduces dynamic datasets along with motion supervision. We found that static datasets play a crucial role in stabilizing training for dynamic scenes, without which the loss would be highly unstable. Since modeling dynamic scenes requires reconstructing a set of 3DGS for each query time, which results in high GPU memory usage, we start by training on 5 input views for dynamic scenes and then expand to 13 views for fine-tuning. Finally, the training resolution is increased to $518\times 518$.
@@ A.3 Training Details
Image encoder, feature backbone and depth head of MoVieS are initialized from a geometrically pretrained transformer, VGGT [10]. Motion head is initialized from its point head. Remaining components, such as the splatter head and camera/time embeddings, are trained from scratch. We use AdamW optimizer [92] with a weight decay of 0.05, and adopt a cosine learning rate scheduler [102] with linear warm-up for all curriculum training stages. For static pretraining and dynamic scenes with 5 and 13 input views at a resolution of $224\times 224$, we use learning rates of 4e-4, 4e-4 and 4e-5, respectively, with a batch size of 256. Training rates for parameters initialized from VGGT are multiplied by 0.1. With 32 H20 GPUs, training takes about 2 days for static and then dynamic scenes with 5 views, and 2 days for 13 views. MoVieS is then finetuned on $518\times 518$ videos with 13 frames using a learning rate of 1e-5, which takes around 1 day. To improve memory and computation efficiency, several techniques such as gsplat [93] rendering backend, DeepSpeed [94], gradient checkpointing [95], gradient accumulation, and bf16 mixed precision are adopted in this work.
Appendix B Discussion
B.1 Limitations and Future Work
@@
Moreover, scaling MoVieS to handle large-scale scenes and achieve high-resolution rendering remains a challenge. The computational cost and memory demand grow significantly with scene complexity and resolution, which constrains practical deployment. Developing more compact and efficient dynamic scene representations, possibly through novel model architectures or sparse encoding strategies, is essential to push 4D reconstruction towards real-world applications.
@@
B.2 Broader Impact
MoVieS provides substantial advantages for fields including robotics simulation, AR/VR, autonomous driving, and digital twins by enabling fast and generalizable dynamic scene understanding. Its ability to reconstruct dynamic environments efficiently can accelerate innovation and improve system performance in these areas. However, such powerful technology also raises risks related to the unauthorized generation of content and potential privacy violations. To mitigate these risks, it is essential to establish clear ethical guidelines and implement appropriate regulatory measures to ensure the responsible and safe use of this technology.
Appendix C More Visualization Results
We provide more visualization results of MoVieS in Figure 6, 7, 8 and 9.
Appendix D License Information
@@
Several public datasets from diverse sources are utilized to train our models, including: (1) RealEstate10K [28] [^8] https://google.github.io/realestate10k/ (Creative Commons Attribution 4.0 International License), (2) TartanAir [29] [^9] https://theairlab.org/tartanair-dataset/ (Creative Commons Attribution 4.0 International License), (3) MatrixCity [30] [^10] https://city-super.github.io/matrixcity/ (Apache License), (4) PointOdyssey [33] [^11] https://pointodyssey.com/ (MIT License), (5) DynamicReplica [34] [^12] https://github.com/facebookresearch/dynamic_stereo (Attribution-NonCommercial 4.0 International License), (6) Spring [31] [^13] https://spring-benchmark.org/ (CC BY 4.0), (7) VKITTI2 [32] [^14] https://europe.naverlabs.com/research/computer-vision/proxy-virtual-worlds-vkitti-2/ (Creative Commons Attribution-NonCommercial-ShareAlike 3.0), and (8) Stereo4D [35] [^15] https://stereo4d.github.io/ (CC0 1.0 Universal).
Figure 6: Qualitative results for reconstruction. “Motion” means the 3D points’ movements of the input frame with respect to the first one.
Figure 7: Qualitative results for 3D point tracking. “Motion” means the 3D points’ movements of the input frame with respect to the first one.
Figure 8: More results for scene flow estimation. “Flow” means the optical flow of the input pixels with respect to the first frame. Arrows on pictures roughly mark directions.
Figure 9: More results for moving object segmentation. Masks for moving parts, which are filtered from motion maps, are highlighted across frames in light red.

### v2

@@ 4.4 Ablation and Analysis
We investigate different camera conditioning strategies in the static pretraining stage. As shown in Table 5, camera tokens are injected throughout the feature backbone, enabling effective camera-aware modeling. In contrast, Plücker embeddings provide limited conditioning on their own and are merely comparable to having no camera information at all. However, as a pixel-aligned representation, Plücker embeddings are complementary to camera tokens, and their combination yields the most effective camera conditioning.
@@
To learn 3D movements in dynamic scenes, we provide two kinds of motion supervision: (1) point-wise L1 loss and (2) distribution loss, as described in Equation 5. Their efficiency is evaluated via the 3D point tracking task in Table 5. Without any motion supervision, i.e., learning solely from novel view synthesis, training exhibits severe loss oscillations and frequent None gradients. The distribution loss captures only relative motion between pixels, while the point-wise L1 loss produces more reasonable motion maps. Combining both leads to sharper boundaries. Qualitative results of estimated motions from different motion objectives are provided in Figure 4.
@@
Thanks to the unified design of MoVieS, it supports simultaneous novel view synthesis (NVS) and motion estimation. We study their mutual benefits in Table 6. “NVS w/o motion” disables explicit motion supervision during training, relying solely on NVS as a proxy to learn dynamics. As shown in Table 6 and Figure 4, this setting fails to learn meaningful motion and tends to model static scenes. “Motion w/o NVS” detaches the motion head from 3DGS rendering and instead conditions the depth head on time. Although explicit supervision enables some motion learning, the predictions are blurry and low-quality, as shown in Figure 4. Moreover, the depth head must now model both geometry and dynamics, increasing its burden and negatively affecting NVS. These results highlight the mutual reinforcement between NVS and motion estimation, where joint training leads to better performance on both.
VGGT Initialization
Although MoVieS is initialized by a pretrained model, i.e., VGGT [68], it is not an essential component of the proposed method. We verified it in our preliminary experiments that using a feature backbone trained entirely from scratch still leads to comparable performance with about 3 times slower convergence. The pretrained VGGT backbone simply serves to accelerate training, but does not fundamentally change the outcome. Specifically, we conduct quantitative experiments on DyCheck at $224\times 224$ resolution. MoVieS initialized from VGGT reaches 17.97 dB mPSNR within 2 days, whereas the counterpart trained from scratch achieves 16.64 dB after 6 days and has not yet fully converged, suggesting that continuous training could further improve its performance.
@@ 4.5 Zero-shot Applications
Scene flow can be naturally derived by transforming the estimated per-pixel motion vectors from world coordinates to the target camera’s coordinates. Visualization results in Figure 5(a) present sharp edges and accurate motion directions. Different colors in the flow maps indicate different movement directions that are marked by block arrows. More visualization results are provided in the supplementary material.
@@
By thresholding the norm of per-pixel motion vectors, the estimated motion maps can also be used to segment moving objects (Figure 5(b)), which is an essential task in computer vision and robotics [78, 18]. Remarkably, this is achieved without any explicit supervision during training, demonstrating the strong potential of our method. More visualization results are provided in the supplementary material.
@@ 5 Conclusion
We introduce MoVieS, a feed-forward model for dynamic novel view synthesis from monocular videos. It’s trained on large-scale, diverse datasets and jointly models scene appearance, geometry, and motion in a unified and efficient framework. Dynamic splatter pixels are proposed to represent dynamic scenes, enabling accurate and temporally coherent 4D reconstruction. MoVieS also supports various applications, such as depth estimation, 3D point tracking, scene flow estimation, and moving object segmentation, showcasing its versatility for dynamic scene perception. We hope this work could serve as a step toward generalizable dynamic scene understanding, and broader applications requiring spatial intelligence. Limitations and future work are discussed in the supplementary material.
Supplementary Material
@@ A.1 Dataset Details
Table S.1: Training Datasets. Eight datasets from diverse sources are utilized to train MoVieS at scale. “#Repeat” denotes dataset duplication count during integration to balance their contributions.
Dataset Dynamic? Depth? Tracking? Real? #Scenes #Frames #Repeat
[…]
Detailed information about the training datasets used in this work is provided in Table S.1. The majority of training datasets used in our experiments are synthetic, providing rich annotations such as pixel-aligned depth maps and accurate camera intrinsics and extrinsics. Ground-truth 3D point trajectories are also available in PointOdyssey [96] and DynamicReplica [21]. On the other hand, Stereo4D [20] is a large-scale dataset constructed from YouTube stereo videos and annotated using pretrained foundation models [64, 10]. Despite being partially noisy, its diversity and scale offer strong generalization benefits.
The pretrained backbone, VGGT [68], requires all 3D scenes to be normalized to a unit scale on average, which in turn necessitates depth maps and camera extrinsics to be defined in a consistent metric space. It is satisfied by all synthetic datasets and Stereo4D, whose camera poses and depths are metric-aligned by construction. However, RealEstate10K [97] only provides relative camera parameters estimated via COLMAP [56], resulting in an unknown global scale. To address this issue, we re-estimate both depth maps and camera extrinsics using recent foundation models, including Video Depth Anything [6], Depth Pro [2] and MegaSaM [31], to recover aligned geometry across frames.
@@ A.2 Curriculum Training
We observed that training MoVieS is particularly unstable: the training loss often fluctuates abruptly, and gradients are prone to becoming None. It may arise from sparse annotations and the heterogeneous nature of training datasets, which mix datasets from various sources with differing domains (e.g., indoor vs. outdoor, real vs. synthetic), camera FoV, recording frame rates, etc.
Thanks to the versatile design of MoVieS, we employ a curriculum strategy that gradually increases the training complexity. It begins by pretraining the model on low-resolution ($224\times 224$) static datasets with only depth and photometric losses, then introduces dynamic datasets along with motion supervision. We found that static datasets play a crucial role in stabilizing training for dynamic scenes, without which the loss would be highly unstable. Since modeling dynamic scenes requires reconstructing a set of 3DGS for each query time, which results in high GPU memory usage, we start by training on 5 input views for dynamic scenes and then expand to 13 views for fine-tuning. Finally, the training resolution is increased to $518$. Similar to VGGT [68], in the last training stage, we randomly sample the frame number from $2\sim 13$ and the aspect ratio from $0.5\sim 2$, with the largest side fixed to $518$.
@@ A.3 Training Details
Image encoder, feature backbone, and depth head of MoVieS are initialized from VGGT [68]. Motion head is initialized from its pointmap head. Remaining components, such as the splatter head and camera/time embeddings, are trained from scratch. We use AdamW optimizer [41] with a weight decay of 0.05, and adopt a cosine learning rate scheduler [40] with linear warm-up for all curriculum training stages. For static pretraining and dynamic scenes with 5 and 13 input views at a resolution of $224\times 224$, we use learning rates of 4e-4, 4e-4 and 4e-5, respectively, with a batch size of 256. Training rates for parameters initialized from VGGT are multiplied by 0.1. With 32 H20 GPUs, training takes about 2 days for static and then dynamic scenes with 5 views, and 2 days for 13 views. MoVieS is then finetuned on $518\times 518$ videos with 13 frames using a learning rate of 1e-5, which takes around 1 day. To improve memory and computation efficiency, several techniques such as gsplat [89] rendering backend, DeepSpeed [55], gradient checkpointing [7], gradient accumulation, and bf16 mixed precision are also adopted.
For the multi-task training objective, we did not manually tune the relative weights between different loss terms. Instead, the weights were set to bring the numerical values of the losses into roughly similar ranges: $\lambda_{\text{d}}=\lambda_{\text{r}}=\lambda_{\text{pt}}=1$ and $\lambda_{\text{m}}=\lambda_{\text{dist}}=10$.
Appendix B Limitations and Future Work
@@
Moreover, scaling MoVieS to handle long videos and achieve high-resolution rendering remains a challenge. The computational cost and memory demand grow significantly with scene complexity and resolution, which constrains practical deployment. Developing more compact and efficient dynamic scene representations, possibly through novel model architectures or sparse encoding strategies, is essential to push 4D reconstruction towards real-world applications.
@@
Appendix C License Information
@@
Datasets from diverse sources are utilized to train MoVieS, including: (1) RealEstate10K [97] [^8] https://google.github.io/realestate10k/ (Creative Commons Attribution 4.0 International License), (2) TartanAir [75] [^9] https://theairlab.org/tartanair-dataset/ (Creative Commons Attribution 4.0 International License), (3) MatrixCity [29] [^10] https://city-super.github.io/matrixcity/ (Apache License), (4) PointOdyssey [96] [^11] https://pointodyssey.com/ (MIT License), (5) DynamicReplica [21] [^12] https://github.com/facebookresearch/dynamic_stereo (Attribution-NonCommercial 4.0 International License), (6) Spring [44] [^13] https://spring-benchmark.org/ (CC BY 4.0), (7) VKITTI2 [3] [^14] https://europe.naverlabs.com/research/computer-vision/proxy-virtual-worlds-vkitti-2/ (Creative Commons Attribution-NonCommercial-ShareAlike 3.0), and (8) Stereo4D [20] [^15] https://stereo4d.github.io/ (CC0 1.0 Universal).
Appendix D Broader Impact
MoVieS provides substantial advantages for fields including robotics simulation, AR/VR, autonomous driving, and digital twins by enabling fast and generalizable dynamic scene understanding. Its ability to efficiently reconstruct dynamic environments can accelerate innovation and improve system performance in these areas. However, such powerful technology also raises risks related to the unauthorized generation of content and potential privacy violations. To mitigate these risks, it is essential to establish clear ethical guidelines and implement appropriate regulatory measures to ensure responsible and safe usage.
Appendix E More Visualization Results
We provide more visualization results on novel view synthesis, 3D point tracking, scene flow estimation, and dynamic object segmentation in Figure S.1, S.2, S.3 and S.4 respectively. Qualitative visualizations of novel view synthesis and 3D point tracking are conducted on the DAVIS dataset [49] and TAPVid-3D [25] respectively, which are not included in the training datasets. Other visualization results are conducted on the test set of the curated datasets to evaluate the generalizability of MoVieS. XYZ values in the 3D space of motion maps are normalized to $[0,1]$ and treated as RGB channels for visualization.
Figure S.1: Qualitative results of novel view synthesis on the DAVIS dataset [49]. Camera poses are estimated by MegaSaM [31]. “Motion” denotes the 3D pixel displacements between the “Target Time” frame with respect to the “Target Viewpoint” frame. The “Novel View Synthesis” results show the reconstructed scene at the target time from the target viewpoint.
Figure S.2: Qualitative results for 3D point tracking. “Motion” means the 3D points’ movements of the input frame with respect to the first one. Points in the 3D space are projected to the image space for 2D point tracking visualization.
Figure S.3: Qualitative results for scene flow estimation. “Flow” means the optical flow of the input pixels with respect to the first frame, and is obtained by projecting the estimated motion maps in the 3D space to the camera space. Each color of optical flow means a specific 2D direction and arrows on pictures roughly mark the directions.
Figure S.4: Qualitative results for moving object segmentation. Masks for moving parts, which are filtered from the values of estimated motion maps, are highlighted across frames in light red. Regions with high motion values are regarded as moving parts.